OpenStudy (anonymous):

the difference quotient of f(x)=4x^2-x+2

OpenStudy (anonymous):

Difference quotient: \(\dfrac{f(x+h)-f(x)}{h}\) If \(f(x)=4x^2-x+2\), then \[\begin{align*}f(x+h)&=4(x+h)^2-(x+h)+2\\ &=4\left(x^2+2xh+h^2\right)-x-h+2\\ &=4x^2+8xh+4h^2-x-h+2 \end{align*}\] So you have \[\begin{align*}\frac{f(x+h)-f(x)}{h}&=\frac{\left(4x^2+8xh+4h^2-x-h+2\right)-\left(4x^2-x+2\right)}{h}\\ &=\frac{8xh+4h^2-h}{h}\\ &=8x+4h-1 \end{align*}\]

OpenStudy (anonymous):

thanks